# Quantum Pseudorandomness and Classical Complexity

### v4

@@ 2.3 Haar measure and concentration
We use $\mathbb{S}(N)$ to denote the set of $N$-dimensional pure quantum states, and $\mathbb{U}(N)$ to denote the group of $N\times N$ unitary matrices. When $N=2^{n}$, we identify these with $n$-qubit states and unitary transformations, respectively. We use $\sigma_{N}$ to denote the Haar measure on $\mathbb{S}(N)$, and we let $\mu_{N}$ denote the Haar measure over $\mathbb{U}(N)$. We write $\mathbb{U}(N)^{M}$ for the space of $MN\times MN$ block-diagonal unitary matrices, where each block has size $N\times N$, and we also identify $\mathbb{U}(N)^{M}$ with $M$-tuples of $N\times N$ unitary matrices. We use $\mu_{N}^{M}$ to denote the product measure $\mu_{N}^{M}(U_{1},U_{2},\ldots,U_{M})=\mu_{N}(U_{1})\cdot\mu_{N}(U_{2})\cdots\mu_{N}(U_{M})$ on $\mathbb{U}(N)^{M}$.
@@ 2.4 Complexity theory
A promise problem $\Pi:\{0,1\}^{*}\to\{0,1,\bot\}$ is in $\mathsf{PromisePostBQP}$ (Postselected Bounded-error Quantum Polynomial time) if there exists a randomized polynomial-time classical algorithm $\mathcal{A}(x)$ that outputs a trit $\{0,1,*\}$ such that:
@@
Technically, the definition of $\mathsf{PostBQP}$ is sensitive to the choice of universal gate set used to specify quantum algorithms, as was observed by Kuperberg [Kup15]. However, for most “reasonable” gate sets, such as unitary gates with algebraic entries [Kup15], the choice of gate set is irrelevant. We assume such a gate set, e.g. $\{\mathrm{CNOT},H,T\}$.
We require the following equivalent characterization of $\mathsf{PostBQP}$:
@@ 2.5 Quantum oracles
We frequently consider quantum algorithms that query quantum oracles. In this work, unless otherwise specified, we define queries to a unitary matrix $\mathcal{U}$ to mean a single application of either $\mathcal{U}$ or controlled-$\mathcal{U}$. For technical reasons, we do not allow queries to $\mathcal{U}^{\dagger}$, although we believe that all of the results proved in this work should still hold if queries to $\mathcal{U}^{\dagger}$ are allowed. [^7] Indeed, the only step that does not seem to remain true with inverses is in the proof of Lemma 22 below. We use superscript notation for algorithms that query oracles. For instance, $\mathcal{A}^{\mathcal{U}}(x,|\psi\rangle)$ denotes a quantum algorithm $\mathcal{A}$ that queries an oracle $\mathcal{U}$ and receives a classical input $x$ and a quantum input $|\psi\rangle$.
We consider versions of $\mathsf{PromiseBQP}$, $\mathsf{PromiseQMA}$, and $\mathsf{PromisePostBQP}$ augmented with quantum oracles, where the algorithm (or in the case of $\mathsf{PromiseQMA}$, the verifier) can apply unitary transformations from an infinite sequence $\mathcal{U}=\{\mathcal{U}_{n}\}_{n\in\mathbb{N}}$. We denote the respective complexity classes by $\mathsf{PromiseBQP}^{\mathcal{U}}$, $\mathsf{PromiseQMA}^{\mathcal{U}}$, and $\mathsf{PromisePostBQP}^{\mathcal{U}}$. We assume the algorithm incurs a cost of $n$ to query $\mathcal{U}_{n}$ so that a polynomial-time algorithm on input $x$ can query $\mathcal{U}_{n}$ for any $n\leq\mathrm{poly}(|x|)$. In this model, a query to $\mathcal{U}_{n}$ consists of a single application of either $\mathcal{U}_{n}$ or controlled-$\mathcal{U}_{n}$.
@@ 2.6 Cryptography
A family of functions $\{f_{k}\}_{k\in\{0,1\}^{\kappa}}$ where $f_{k}:\{0,1\}^{n}\to\{0,1\}^{m}$ is called $t$-wise independent if for every distinct $x_{1},x_{2},\ldots,x_{t}\in\{0,1\}^{n}$ and every (not necessarily distinct) $y_{1},y_{2},\ldots,y_{t}\in\{0,1\}^{m}$:
$\Pr_{k\sim\{0,1\}^{\kappa}}\left[f_{k}(x_{i})=y_{i}\;\forall i\in[t]\right]=2^{-mt}.$
Efficient constructions of $t$-wise independent functions are known, in the sense that one can sample a random $f_{k}$ from a $t$-wise independent function family and make queries to $f_{k}$ in $\mathrm{poly}(t,n,m)$ time [Zha12]. Our primary use of $t$-wise independent functions is in simulating random oracles: $2t$-wise independent functions can be used in place of a uniformly random function $\{0,1\}^{n}\to\{0,1\}^{m}$ in any quantum algorithm that makes at most $t$ queries to the random function; see Zhandry [Zha12, Theorem 6.1] for more details.
@@
    (Computationally indistinguishable) For any polynomial-time quantum adversary $\mathcal{A}$ and every $T=\mathrm{poly}(\kappa)$:
@@
We sometimes call the negligible quantities in the above definitions the advantage of the quantum adversary $\mathcal{A}$.
In this work, we consider security against non-uniform quantum algorithms with classical advice, which means that the adversary is allowed to be a different polynomial-time quantum algorithm for each setting of the security parameter $\kappa\in\mathbb{N}$. Without loss of generality, such an adversary can always be assumed to take the form of a uniform $\mathrm{poly}(\kappa)$-time quantum algorithm $\mathcal{A}\left(1^{\kappa},x\right)$, where $x\in\{0,1\}^{\mathrm{poly}(\kappa)}$ is an advice string that depends only on $\kappa$.
@@ 3 Approximate $t$-designs
For each $n,t\in\mathbb{N}$ and $\varepsilon>0$, there exists $m(n)\leq\mathrm{poly}(n,t,\log\frac{1}{\varepsilon})$ and a $\mathrm{poly}(n,t,\log\frac{1}{\varepsilon})$-time classical algorithm $\mathcal{S}$ that takes as input a random string $x\sim\{0,1\}^{m}$ and outputs a description of a quantum circuit on $n$ qubits such that the circuits sampled from $\mathcal{S}$ form an $\varepsilon$-approximate unitary $t$-design over $\mathbb{U}(2^{n})$.
@@
$(1-\varepsilon)\Pr_{U\sim\mu_{N}}\left[\mathcal{A}^{U}=1\right]\leq\Pr_{U\sim S}\left[\mathcal{A}^{U}=1\right]\leq(1+\varepsilon)\Pr_{U\sim\mu_{N}}\left[\mathcal{A}=1\right].$
@@
where $|\Phi_{N}\rangle=\sum_{i=1}^{N}|i\rangle|i\rangle$.
@@
Let $\mathcal{A}^{U}$ be a quantum algorithm that makes $t$ adaptive queries to $U\in\mathbb{U}(N)$ (but not controlled-$U$). Then there exists an algorithm $\mathcal{B}\left(|\phi_{U}\rangle^{\otimes t}\right)$ such that:
@@
Let $\mathcal{B}$ simulate each query that $\mathcal{A}$ makes to $U$ using the above circuit and a copy of $|\phi_{U}\rangle$. Conditioned on all $t$ postselection steps succeeding, which occurs with probability $\frac{1}{N^{2t}}$, the output state of $\mathcal{B}$ is exactly the same as the the output state of $\mathcal{A}$. Thus,
@@
An interesting question is whether Lemma 22 can be generalized to algorithms $\mathcal{A}^{U}$ that can make queries to both $U$ and $U^{\dagger}$. If true, this would imply that all further results proved in this work could be extended to algorithms that make queries to both $U$ and $U^{\dagger}$. In a sense, we are just asking whether queries to $U^{\dagger}$ can be simulated by a combination of queries to $U$ and postselection. Notably, in the case that $U$ is a real orthogonal matrix, this is possible, because $|\phi_{U}\rangle$ is equivalent to $|\phi_{U^{\top}}\rangle$ up to swapping the two registers, and $U^{\top}=U^{\dagger}$ for real matrices $U$.
@@
Let $S$ be an $\varepsilon$-approximate unitary $t$-design over $\mathbb{U}(N)$, and let $\mathcal{A}^{U}$ be a quantum algorithm that makes $t$ queries to $U$ (but not controlled-$U$). Then:
$(1-\varepsilon)\Pr_{U\sim\mu_{N}}\left[\mathcal{A}^{U}=1\right]\leq\Pr_{U\sim S}\left[\mathcal{A}^{U}=1\right]\leq(1+\varepsilon)\Pr_{U\sim\mu_{N}}\left[\mathcal{A}=1\right].$
@@
Finally, we wish to extend Lemma 23 to algorithms that can also make queries to controlled-$U$. It is not at all obvious that this is possible, because it is well known that queries to controlled-$U$ cannot be simulated efficiently using queries to $U$ [GST20]. At the same time, it is unclear how one should pick the phase on controlled-$U$, because the definition of an approximate unitary $t$-design (Definition 17) “forgets” the global phase of $U$. We address both problems simultaneously by observing that the argument using Lemma 22 can still go through if we choose the phase of $U$ randomly. In order for this argument to hold, we assume that our approximate unitary $t$-design $S$ is phase-invariant, which we take to mean that for any $U$ sampled from the design, $U$ and $\omega U$ are chosen with the same probability, where $\omega=e^{\frac{2\pi i}{t+1}}$ is a primitive $(t+1)$th root of unity. [^8] Actually, our proof only requires that the design be invariant under any group of complex units whose first $t$ moments are the same as the uniform measure over complex units. We choose the group generated by $\omega$ only because it is the smallest group with this property. Note that we can make any unitary design phase-invariant via multiplication by a uniformly random $(t+1)$th root of unity.
@@
Let $S$ be a phase-invariant $\varepsilon$-approximate unitary $t$-design over $\mathbb{U}(N)$, and let $\mathcal{A}^{U}$ be a quantum algorithm that makes $t$ queries to $U$ (including controlled-$U$). Then:
$(1-\varepsilon)\Pr_{U\sim\mu_{N}}\left[\mathcal{A}^{U}=1\right]\leq\Pr_{U\sim S}\left[\mathcal{A}^{U}=1\right]\leq(1+\varepsilon)\Pr_{U\sim\mu_{N}}\left[\mathcal{A}=1\right].$
@@
Let $\langle\omega\rangle$ be the group of $(t+1)$th roots of unity. We first claim that given $t$ copies of the Choi state of $U$ (i.e. $|\phi_{U}\rangle^{\otimes t}$), we can efficiently generate the state
@@
By identifying $[2N]$ with $\{0,1\}\times[N]$, observe that
$|\phi_{I\oplus\varphi U}\rangle=\frac{|0\rangle|\Phi_{N}\rangle+\varphi|1\rangle|\phi_{U}\rangle}{\sqrt{2}}$
For convenience, define $|\psi_{0}\rangle\coloneqq|0\rangle|\Phi_{N}\rangle$ and $|\psi_{1}\rangle\coloneqq|1\rangle|\phi_{U}\rangle$. For $x\in\{0,1\}^{t}$, we extend this notation via $|\psi_{x}\rangle\coloneqq\bigotimes_{i=1}^{t}|\psi_{x_{i}}\rangle$. For $i\in[t]$, define
@@
$\rho_{U,t}\coloneqq\sum_{i\in[t]}\frac{\binom{t}{i}}{2^{t}}|\overline{\psi}_{i}\rangle\langle\overline{\psi}_{i}|.$
@@ 4 Breaking pseudorandomness with a classical oracle
Let $|\varphi\rangle\in\mathbb{S}(N)$, and let $\varepsilon>0$. Then: [^9] As observed in [${}^{+}$23], an earlier version of this work [Kre21a] gave the incorrect bound $e^{-\varepsilon N}$ here, which was carried over from [BHH16b, Equation (14)].
@@
For any PRS ensemble $\{|\varphi_{k}\rangle\}_{k\in\{0,1\}^{\kappa}}$ of $n$-qubit states with security parameter $\kappa$, there exists a $\mathsf{PP}$ language $\mathcal{L}$, a $\mathrm{poly}(\kappa)$-time quantum algorithm $\mathcal{A}^{\mathcal{L}}$, and $T=\mathrm{poly}(\kappa)$ such that the following holds. Let $X\sim\{0,1\}$ be a uniform random bit. Let $|\psi\rangle$ be sampled uniformly from the PRS ensemble if $X=0$, and otherwise let $|\psi\rangle$ be sampled from the Haar measure $\sigma_{2^{n}}$ if $X=1$. Then we have:
@@
The $\mathsf{PP}$ language $\mathcal{L}$ that we choose for the oracle is most easily described in terms of a $\mathsf{PromisePostBQP}$ algorithm $\mathcal{B}(b,c)$ (i.e. a postselected polynomial-time quantum algorithm, as in Definition 12), by the equivalence $\mathsf{PostBQP}=\mathsf{PP}$ (Lemma 13). [^10] Technically, the $\mathsf{PromisePostBQP}$ algorithm we define has a promise, so a priori it only defines a promise problem and not a language. However, by Aaronson’s theorem (Lemma 13) any promise problem in $\mathsf{PromisePostBQP}$ is also in $\mathsf{PromisePP}$ [Aar05], and any promise problem in $\mathsf{PromisePP}$ can be extended to a language in $\mathsf{PP}$ because $\mathsf{PP}$ is a syntactic class. Hence, we might as well take a language in $\mathsf{PP}$ instead of a promise problem. Let $S$ be a $\frac{1}{17}$-approximate $n$-qubit quantum $T$-design (Definition 16) such that a state can be drawn from $S$ in $\mathrm{poly}(\kappa)$ time (because $n,T\leq\mathrm{poly}(\kappa)$, the existence of such a design follows from Lemma 19). $\mathcal{B}$ begins by initializing the state:
$\hat{\rho}=\frac{1}{2}|0\rangle\langle 0|\otimes\mathop{\mathbb{E}}_{k\sim\{0,1\}^{\kappa}}\left[|\varphi_{k}\rangle\langle\varphi_{k}|^{\otimes T}\right]+\frac{1}{2}|1\rangle\langle 1|\otimes\mathop{\mathbb{E}}_{|\psi\rangle\sim S}\left[|\psi\rangle\langle\psi|^{\otimes T}\right].$
$\mathcal{B}$ measures all but the leftmost qubit of $\hat{\rho}$ in the basis given by $b$, and postselects on observing $c$ (i.e. $\mathcal{B}$ outputs *** if the measurements are not equal to $c$). Finally, conditioned on postselection succeeding, $\mathcal{B}$ measures and outputs the result of the leftmost qubit that was not measured.
@@
$\rho=\frac{1}{2}|0\rangle\langle 0|\otimes\mathop{\mathbb{E}}_{k\sim\{0,1\}^{\kappa}}\left[|\varphi_{k}\rangle\langle\varphi_{k}|^{\otimes T}\right]+\frac{1}{2}|1\rangle\langle 1|\otimes\mathop{\mathbb{E}}_{|\psi\rangle\sim\sigma_{2^{n}}}\left[|\psi\rangle\langle\psi|^{\otimes T}\right],$

### v5

@@ 2.4 Haar measure and concentration
We use $\mathbb{S}(N)$ to denote the set of $N$-dimensional pure quantum states, and $\mathbb{U}(N)$ to denote the group of $N\times N$ unitary matrices. When $N=2^{n}$, we identify these with $n$-qubit states and unitary transformations, respectively. We use $\sigma_{N}$ to denote the Haar measure on $\mathbb{S}(N)$, and we let $\mu_{N}$ denote the Haar measure over $\mathbb{U}(N)$. We write $\mathbb{U}(N)^{M}$ for the space of $MN\times MN$ block-diagonal unitary matrices, where each block has size $N\times N$, and we also identify $\mathbb{U}(N)^{M}$ with $M$-tuples of $N\times N$ unitary matrices. We use $\mu_{N}^{M}$ to denote the product measure $\mu_{N}^{M}(U_{1},U_{2},\ldots,U_{M})\coloneqq\mu_{N}(U_{1})\cdot\mu_{N}(U_{2})\cdots\mu_{N}(U_{M})$ on $\mathbb{U}(N)^{M}$, which we interpret as a distribution over a direct sum $U_{1}\oplus U_{2}\oplus\ldots\oplus U_{M}$ of matrices.
@@ 2.5 Complexity theory
$\mathsf{QMA}$ is defined as the set of languages in $\mathsf{PromiseQMA}$.
Note: we will sometimes call any algorithm of the form $\mathcal{V}(x,|\psi\rangle)$ a $\mathsf{QMA}$ verifier, even if it does not satisfy the promise of a $\mathsf{QMA}$ language.
@@
A promise problem $\Pi:\{0,1\}^{*}\to\{0,1,\bot\}$ is in $\mathsf{PromisePostBQP}$ (Postselected Bounded-error Quantum Polynomial time) if there exists a polynomial-time quantum algorithm $\mathcal{A}(x)$ that outputs a trit in $\{0,1,*\}$ such that:
@@
$\mathsf{PostBQP}$ is defined as the set of languages in $\mathsf{PromisePostBQP}$.
Technically, the definition of $\mathsf{PromisePostBQP}$ is sensitive to the choice of universal gate set used to specify quantum algorithms, as was observed by Kuperberg [Kup15]. However, for most “reasonable” gate sets, such as unitary gates with algebraic entries [Kup15], the choice of gate set is irrelevant. We assume such a gate set, e.g. $\{\mathrm{CNOT},H,T\}$.
We require the following equivalent characterization of $\mathsf{PromisePostBQP}$:
@@ 2.6 Quantum oracles
We frequently consider quantum algorithms that query quantum oracles. In this work, unless otherwise specified, we define queries to a unitary matrix $\mathcal{U}$ to mean a single application of either $\mathcal{U}$, $\mathcal{U}^{\dagger}$, controlled-$\mathcal{U}$ (i.e. $I\oplus\mathcal{U}$, where $I$ is the identity of the same dimension), or controlled-$\mathcal{U}^{\dagger}$ (i.e. $I\oplus\mathcal{U}^{\dagger}$), unless otherwise specified. We use superscript notation for algorithms that query oracles. For instance, $\mathcal{A}^{\mathcal{U}}(x,|\psi\rangle)$ denotes a quantum algorithm $\mathcal{A}$ that queries an oracle $\mathcal{U}$ and receives a classical input $x$ and a quantum input $|\psi\rangle$.
We consider versions of $\mathsf{PromiseBQP}$, $\mathsf{PromiseQMA}$, and $\mathsf{PromisePostBQP}$ augmented with quantum oracles, where the algorithm (or in the case of $\mathsf{PromiseQMA}$, the verifier) can apply unitary transformations from an infinite sequence $\mathcal{U}=\{\mathcal{U}_{n}\}_{n\in\mathbb{N}}$. We denote the respective complexity classes by $\mathsf{PromiseBQP}^{\mathcal{U}}$, $\mathsf{PromiseQMA}^{\mathcal{U}}$, and $\mathsf{PromisePostBQP}^{\mathcal{U}}$. We assume the algorithm incurs a cost of $n$ to query $\mathcal{U}_{n}$ so that a polynomial-time algorithm on input $x$ can query $\mathcal{U}_{n}$ for any $n\leq\mathrm{poly}(|x|)$. In this model, a query to $\mathcal{U}_{n}$ consists of a single application of either $\mathcal{U}_{n}$, controlled-$\mathcal{U}_{n}$, or their inverses.
@@ 2.7 Cryptography
    (Computationally indistinguishable) For any polynomial-time quantum adversary $\mathcal{A}$ and for every $T=\mathrm{poly}(\kappa)$:
@@
We sometimes call the negligible quantities in the above definitions the advantage of the quantum adversary $\mathcal{A}$. Additionally, we may instantiate these primitives relative to an oracle $\mathcal{O}$, which just means that both the generating algorithm $G$ and the adversary $\mathcal{A}$ are additionally allowed to query $\mathcal{O}$.
In this work, we will only consider pseudorandom state and unitary ensembles where $n(\kappa)=\omega(\log\kappa)$. Although the original definition of Ji, Liu, and Song [JLS18] did not impose this condition, later works have shown that $O(\log\kappa)$-qubit pseudorandom ensembles behave very differently from $\omega(\log\kappa)$-qubit ensembles [BS20, AQY22, BEM24]; the former are often called “short PRSs/PRUs”. Intuitively, this difference is because one can perform tomography on a quantum state or unitary of $O(\log\kappa)$ qubits to any desired precision $\varepsilon$ in time $\mathrm{poly}(\kappa,\varepsilon)$, so short PRSs/PRUs behave more like cryptographic objects with classical output.
We must also be careful about the type of adversary $\mathcal{A}$ considered in Definitions 15 and 16. In this work, we consider security against non-uniform quantum algorithms with classical advice, which means that the adversary is allowed to be a different polynomial-time quantum algorithm for each setting of the security parameter $\kappa\in\mathbb{N}$. Without loss of generality, such an adversary can always be assumed to take the form of a uniform $\mathrm{poly}(\kappa)$-time quantum algorithm $\mathcal{A}\left(1^{\kappa},x\right)$, where $x\in\{0,1\}^{\mathrm{poly}(\kappa)}$ is an advice string that depends only on $\kappa$.
@@ 3 Approximate $t$-designs
For each $n,t\in\mathbb{N}$ and $\varepsilon>0$, there exists $m\leq\mathrm{poly}(n,t,\log\frac{1}{\varepsilon})$ and a $\mathrm{poly}(n,t,\log\frac{1}{\varepsilon})$-time classical algorithm $\mathcal{S}$ that takes as input a random string $x\sim\{0,1\}^{m}$ and outputs a description of a quantum circuit on $n$ qubits such that the circuits sampled from $\mathcal{S}$ form an $\varepsilon$-approximate unitary $t$-design over $\mathbb{U}(2^{n})$.
@@
$(1-\varepsilon)\Pr_{U\sim\mu_{N}}\left[\mathcal{A}^{U}=1\right]\leq\Pr_{U\sim S}\left[\mathcal{A}^{U}=1\right]\leq(1+\varepsilon)\Pr_{U\sim\mu_{N}}\left[\mathcal{A}^{U}=1\right].$
@@
where $|\Phi_{N}\rangle\coloneqq\frac{1}{\sqrt{N}}\sum_{i=1}^{N}|i\rangle|i\rangle$ is the standard maximally entangled state.
@@
Let $\mathcal{A}^{U}$ be a quantum algorithm that makes $t$ adaptive queries to $U\in\mathbb{U}(N)$ (but not controlled-$U$ or $U^{\dagger}$). Then there exists an algorithm $\mathcal{B}\left(|\phi_{U}\rangle^{\otimes t}\right)$ such that:
@@
Let $\mathcal{B}$ simulate each query that $\mathcal{A}$ makes to $U$ using the above circuit and a copy of $|\phi_{U}\rangle$. Conditioned on all $t$ postselection steps succeeding, which occurs with probability $\frac{1}{N^{2t}}$, the output state of $\mathcal{B}$ is exactly the same as the output state of $\mathcal{A}$. Thus,
@@
An interesting question is whether Lemma 23 can be generalized to algorithms $\mathcal{A}^{U}$ that can make queries to both $U$ and $U^{\dagger}$. In a sense, we are just asking whether queries to $U^{\dagger}$ can be simulated by a combination of queries to $U$ and postselection. Notably, in the case that $U$ is a real orthogonal matrix, this is possible, because $|\phi_{U}\rangle$ is equivalent to $|\phi_{U^{\top}}\rangle$ up to swapping the two registers, and $U^{\top}=U^{\dagger}$ for real matrices $U$.
@@
Let $S$ be an $\varepsilon$-approximate unitary $t$-design over $\mathbb{U}(N)$, and let $\mathcal{A}^{U}$ be a quantum algorithm that makes $t$ queries to $U$ (but not controlled-$U$ or $U^{\dagger}$). Then:
$(1-\varepsilon)\Pr_{U\sim\mu_{N}}\left[\mathcal{A}^{U}=1\right]\leq\Pr_{U\sim S}\left[\mathcal{A}^{U}=1\right]\leq(1+\varepsilon)\Pr_{U\sim\mu_{N}}\left[\mathcal{A}^{U}=1\right].$
@@
Finally, we wish to extend Lemma 24 to algorithms that can also make queries to controlled-$U$. It is not at all obvious that this is possible, because it is well known that queries to controlled-$U$ cannot be simulated efficiently using queries to $U$ [GST24]. At the same time, it is unclear how one should pick the phase on controlled-$U$, because the definition of an approximate unitary $t$-design (Definition 18) “forgets” the global phase of $U$. We address both problems simultaneously by observing that the argument using Lemma 23 can still go through if we choose the phase of $U$ randomly. In order for this argument to hold, we assume that our approximate unitary $t$-design $S$ is phase-invariant, which we take to mean that for any $U$ sampled from the design, $U$ and $\omega U$ are chosen with the same probability, where $\omega=e^{\frac{2\pi i}{t+1}}$ is a primitive $(t+1)$th root of unity. [^5] Actually, our proof only requires that the design be invariant under any group of complex units whose first $t$ moments are the same as the uniform measure over complex units. We choose the group generated by $\omega$ only because it is the smallest group with this property. Note that we can make any unitary design phase-invariant via multiplication by a uniformly random $(t+1)$th root of unity.
@@
Let $S$ be a phase-invariant $\varepsilon$-approximate unitary $t$-design over $\mathbb{U}(N)$, and let $\mathcal{A}^{U}$ be a quantum algorithm that makes $t$ queries to $U$ (including controlled-$U$, but not $U^{\dagger}$). Then:
$(1-\varepsilon)\Pr_{U\sim\mu_{N}}\left[\mathcal{A}^{U}=1\right]\leq\Pr_{U\sim S}\left[\mathcal{A}^{U}=1\right]\leq(1+\varepsilon)\Pr_{U\sim\mu_{N}}\left[\mathcal{A}^{U}=1\right].$
@@
Let $\langle\omega\rangle$ be the group of $(t+1)$th roots of unity. We first claim that given $t$ copies of the Choi state of $U$ (i.e. $|\phi_{U}\rangle^{\otimes t}$), we can generate the state
@@
Recall that $|\phi_{I\oplus\varphi U}\rangle$ is defined by
$|\phi_{I\oplus\varphi U}\rangle\coloneqq\frac{1}{\sqrt{2N}}\left(\sum_{i=1}^{N}|i\rangle|i\rangle+\sum_{i=N}^{2N}\varphi U|i\rangle|i\rangle\right).$
By identifying $[2N]$ with $\{0,1\}\times[N]$, we can also write this in the form
$|\phi_{I\oplus\varphi U}\rangle\equiv\frac{1}{\sqrt{2N}}\left(\sum_{i=1}^{N}|0\rangle|i\rangle|0\rangle|i\rangle+\sum_{i=1}^{N}\varphi U|1\rangle|i\rangle|1\rangle|i\rangle\right).$
Swapping the ordering of the second and third registers, we identify this state with
$|\phi_{I\oplus\varphi U}\rangle\equiv\frac{|00\rangle|\Phi_{N}\rangle+\varphi|11\rangle|\phi_{U}\rangle}{\sqrt{2}}.$
For convenience, define $|\psi_{0}\rangle\coloneqq|00\rangle|\Phi_{N}\rangle$ and $|\psi_{1}\rangle\coloneqq|11\rangle|\phi_{U}\rangle$. For $x\in\{0,1\}^{t}$, we extend this notation via $|\psi_{x}\rangle\coloneqq\bigotimes_{i=1}^{t}|\psi_{x_{i}}\rangle$. For $i=0,\ldots,t$, define
@@
$\rho_{U,t}\coloneqq\sum_{i=1}^{t}\frac{\binom{t}{i}}{2^{t}}|\overline{\psi}_{i}\rangle\langle\overline{\psi}_{i}|.$
@@ 4 Breaking pseudorandomness with a classical oracle
Let $|\varphi\rangle\in\mathbb{S}(N)$, and let $\varepsilon>0$. Then: [^6] As observed in [CGG+23], an earlier version of this work [Kre21a] gave the incorrect bound $e^{-\varepsilon N}$ here, which was carried over from [BHH16b, Equation (14)].
@@
For any PRS ensemble $\{|\varphi_{k}\rangle\}_{k\in\{0,1\}^{\kappa}}$ of $n$-qubit states with security parameter $\kappa$ satisfying $n=\omega(\log\kappa)$, there exists a $\mathsf{PP}$ language $\mathcal{L}$, a $\mathrm{poly}(\kappa)$-time quantum algorithm $\mathcal{A}^{\mathcal{L}}$, and $T=\mathrm{poly}(\kappa)$ such that the following holds. Let $X\sim\{0,1\}$ be a uniform random bit. Let $|\psi\rangle$ be sampled uniformly from the PRS ensemble if $X=0$, and otherwise let $|\psi\rangle$ be sampled from the Haar measure $\sigma_{2^{n}}$ if $X=1$. Then we have:
@@
The $\mathsf{PP}$ language $\mathcal{L}$ that we choose is most easily described in terms of a $\mathsf{PromisePostBQP}$ algorithm $\mathcal{B}(b,c)$ (i.e. a postselected polynomial-time quantum algorithm, as in Definition 13), by the equivalence $\mathsf{PromisePostBQP}=\mathsf{PromisePP}$ (Lemma 14). That is, we specify an algorithm $\mathcal{B}(b,c)$ that outputs a trit in $\{0,1,*\}$, and this algorithm defines a promise problem $\Pi\in\mathsf{PromisePostBQP}$ as follows:
  1. (i)
    If $\Pr[\mathcal{B}(b,c)\in\{0,1\}]>0$ and $\Pr[\mathcal{B}(b,c)=1\mid\mathcal{B}(b,c)\in\{0,1\}]\geq\frac{2}{3}$, then $\Pi(b,c)=1$.
  2. (ii)
    If $\Pr[\mathcal{B}(b,c)\in\{0,1\}]>0$ and $\Pr[\mathcal{B}(b,c)=1\mid\mathcal{B}(b,c)\in\{0,1\}]\leq\frac{1}{3}$, then $\Pi(b,c)=0$.
  3. (iii)
    Otherwise, $\Pi(b,c)=\bot$.
By Aaronson’s theorem (Lemma 14), $\Pi$ is also in $\mathsf{PromisePP}$. Because $\mathsf{PP}$ is a syntactic class, the promise problem $\Pi$ can be extended to a language $\mathcal{L}\in\mathsf{PP}$.
Let $S$ be a $\frac{1}{17}$-approximate $n$-qubit quantum $T$-design (Definition 17) such that a state can be drawn from $S$ in $\mathrm{poly}(\kappa)$ time (because $n,T\leq\mathrm{poly}(\kappa)$, the existence of such a design follows from Proposition 19 and Lemma 20). $\mathcal{B}$ begins by initializing the state:
$\hat{\rho}\coloneqq\frac{1}{2}|0\rangle\langle 0|\otimes\mathop{\mathbb{E}}_{k\sim\{0,1\}^{\kappa}}\left[|\varphi_{k}\rangle\langle\varphi_{k}|^{\otimes T}\right]+\frac{1}{2}|1\rangle\langle 1|\otimes\mathop{\mathbb{E}}_{|\phi\rangle\sim S}\left[|\phi\rangle\langle\phi|^{\otimes T}\right].$
$\mathcal{B}$ measures all but the leftmost qubit of $\hat{\rho}$ in the basis given by $b$, and postselects on observing $c$ (i.e. $\mathcal{B}$ outputs * if the measurements are not equal to $c$). Finally, conditioned on postselection succeeding, $\mathcal{B}$ measures and outputs the result of the leftmost qubit that was not measured.
@@
$\rho\coloneqq\frac{1}{2}|0\rangle\langle 0|\otimes\mathop{\mathbb{E}}_{k\sim\{0,1\}^{\kappa}}\left[|\varphi_{k}\rangle\langle\varphi_{k}|^{\otimes T}\right]+\frac{1}{2}|1\rangle\langle 1|\otimes\mathop{\mathbb{E}}_{|\phi\rangle\sim\sigma_{2^{n}}}\left[|\phi\rangle\langle\phi|^{\otimes T}\right],$